OpenStudy (babynini):

How to solve: tan(theta) = 2/(-2 sqroot3)

OpenStudy (babynini):

\[\tan(\theta)=\frac{ 2 }{ -2\sqrt{3} }\]

OpenStudy (babynini):

\[\theta = \tan ^{-1}(\frac{ 2 }{ -2\sqrt{3} })\]

OpenStudy (babynini):

@Nnesha

OpenStudy (babynini):

@jim_thompson5910

Nnesha (nnesha):

can you cancel out 2's ??

OpenStudy (babynini):

multiply it by 2/2?

OpenStudy (babynini):

ooh I see what you're saying, sorry.

OpenStudy (babynini):

err can I?

OpenStudy (babynini):

then it would be \[\frac{ 1 }{ -\sqrt{3} }\]

Nnesha (nnesha):

\[\huge\rm tan(\theta) =\color{red}{-\frac{ 2 }{ 2\sqrt{3}} }\] first of all solve red part :-)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Babynini then it would be \[\frac{ 1 }{ -\sqrt{3} }\] \(\color{blue}{\text{End of Quote}}\) yes but you ca't have radical at the denominator so multiply top and bottom by sqrt{3}

OpenStudy (babynini):

\[\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (babynini):

with a negative in front :P

Nnesha (nnesha):

\[\theta= \tan^{-1} (-\frac{\sqrt{3}}{3} )\] from there do you wanna use a calculator or draw triangle ?

OpenStudy (babynini):

Triangle! Because what I need is it in radical form

OpenStudy (babynini):

(for example 5pi/6)

Nnesha (nnesha):

 to make a triangle we know that it's not going to be in 1st and 3rd quadrant bec tan is positive in 1st and 3rd q